OpenStudy (anonymous):

7v^3 * 10u^3v^5 * 8uv^3

OpenStudy (tkhunny):

What have you tried? Collect like bases and add exponents.

OpenStudy (unklerhaukus):

\[7v^3 \times 10u^3v^5 \times 8uv^3\] \[=7\times10\times8\times v^{3+5+3}\times u^{3+1}\] \[=\]